OpenStudy (anonymous):

7n+6n-5=4n+4 Show work.

OpenStudy (the_fizicx99):

Combine like terms, 7n + 6n = 13n 13n - 5 = 4n + 4 Add both sides 5 13n - 5 = 4n + 4 +5 +5 13n = 4n + 9 Subtract 4n from both sides 13n = 4n +9 -4n -4n 9n = 9 /9 /9 n = 1

OpenStudy (the_fizicx99):

This is a true statement because , 7(1) + 6(1) - 5 = 4(1) + 4 7 + 6 - 5 = 4 + 4 13 - 5 = 8 4 +4 = 8 8 = 8 <-- true